Consider the Following,

C=40+0.6Y .......1

I=20+0.1Y+0.4r..........2

Y=C+I, where C is Consumption and I is Investment and Y is National Income.

Find the reduced form of the equation of C, I and Y


1
Expert's answer
2021-02-22T14:02:24-0500

C=40+0.6Y=40+0.6(C+I)=40+0.6C+0.6I; 0.4C=40+0.6I; C=100+1.5I

I=20+0.1Y+0.4r=20+0.1(C+I)+0.4r=20+0.1C+0.1I+0.4r; then 0.9I=20+0.1C+0.4r; C=9I-200-4r

Then 100+1.5I=9I-200-4r; I=(300+4r)/7.5=40+0.53r

C=100+1.5(40+0.53r)=100+60+0.8r=160+0.8r

Y=40+0.53r+160+0.8r=200+1.33r